# MotiF: Making Text Count in Image Animation with Motion Focal Loss

### v1

@@ A.2 Motion Focal Loss Weight
We investigate the impact of the motion focal loss weight $\lambda$ on TI2V Bench. As shown in Table A2, simply setting $\lambda=1$ achieves the best overall results. Generally, reducing $\lambda$ can improve the overall visual quality of the generated videos but results in lower TI2V Score specially on worse text alignment and object motion.
@@ B.1 Evaluation on VBench-I2V
VBench-I2V [21] is another popular image-to-video (I2V) benchmark, consisting of 356 real-world images and 1,118 image-prompt pairs. Different from TI2V Bench, VBench-I2V uses image captions as text conditions instead of action instructions and especially focus on controlling camera motion through text prompts. We evaluate MotiF alongside two strongest baseline models, DynamiCrafter and Cinemo, as well as the static video baseline, on VBench-I2V. Results are presented in Table A3.
@@ Appendix C Additional Visualization
Figure A1: More qualitative comparison to prior works on TI2V Bench. MotiF can generate videos that align better with the text prompts. More video samples are available in the project website.
Figure A2: Typical failure and challenging cases of MotiF on TI2V Bench. We observe two typical cases that the model fail: 1) the generated videos may have unnatural motion ((a)); 2) the generated videos do not align well with the prompts ((b), (c), (d)). For 2), there are two specific scenarios when following the text is challenging including novel object ((c)) or multiple objects ((d)). We also include more video samples in the project website.
@@ C.1 Comparison with Baseline Models
C.2 Failing Cases Analysis
Although MotiF shows clear advantages over prior work, it is still far from perfect for TI2V generation especially on our proposed challenging benchmark TI2V Bench. As shown in Figure A2, we observe two main types of failure cases. First, sometimes the generated motion is not very natural. Second, the generated video may not follow the text prompt. In the second case, there are two challenging scenarios of TI2V Bench: 1) when the text prompt describes a new object that needs to appear in the scene, the generated video may not be very coherent; 2) when there are multiple objects and the text prompt only refers to one of them, it will be hard for the model to generate precise motion.
We hope MotiF and TI2V Bench will help the research community to tackle this challenging problem. MotiF is generic and complementary to existing techniques for TI2V generation. We believe that improving the motion heatmap accuracy can potentially boost the performance. Moreover, MotiF is potentially applicable to text-to-video generation.
@@ Appendix D Additional Implementation Details
D.1 Scenes in TI2V Bench
TI2V Bench contains a total of 22 diverse scenes, designed to cover a wide range of scenarios. The detailed scenes include: car on the road, balance scale, (multiple) balloons, bird, (multiple) bulbs, butterfly, candle, child in a playground, dog, rubber duck on a pool, fish, flower, golf ball, horse, animal on a meadow, human face, human body, sun, tide, traffic light, tree, and volcano.
D.2 Human Evaluation
We show the human evaluation interface in A3. During the evaluation process, the annotators are required to read the instructions first and then answer the two questions based on the specified criteria.

### v2

@@ A.2 Motion Focal Loss Weight
We investigate the impact of the motion focal loss weight $\lambda$ on TI2V-Bench. As shown in Table A2, simply setting $\lambda=1$ achieves the best overall results. Generally, reducing $\lambda$ can improve the overall visual quality of the generated videos but results in lower TI2V Score specially on worse text alignment and object motion.
@@ B.1 Evaluation on VBench-I2V
VBench-I2V [21] is another popular image-to-video (I2V) benchmark, consisting of 356 real-world images and 1,118 image-prompt pairs. Different from TI2V-Bench, VBench-I2V uses image captions as text conditions instead of action instructions and especially focus on controlling camera motion through text prompts. We evaluate MotiF alongside two strongest baseline models, DynamiCrafter and Cinemo, as well as the static video baseline, on VBench-I2V. Results are presented in Table A3.
@@ Appendix C Additional Visualization
Figure A1: More qualitative comparison to prior works on TI2V-Bench. MotiF can generate videos that align better with the text prompts. More video samples are available in the project website.
Figure A2: Results on complex scenarios. MotiF generates faithful videos for (1) object occlusion and (b) multiple object interaction.
Figure A3: Typical failure and challenging cases of MotiF on TI2V-Bench. We observe two typical cases that the model fail: 1) the generated videos may have unnatural motion ((a)); 2) the generated videos do not align well with the prompts ((b), (c), (d)). For 2), there are two specific scenarios when following the text is challenging including novel object ((c)) or multiple objects ((d)). We also include more video samples in the project website.
@@ C.1 Comparison with Baseline Models
C.2 Additional Examples
We also provide additional examples that goes beyond TI2V-Bench for complex scenarios including occlusions and multi-object interaction. The results are shown in Figure A2.
C.3 Failing Cases Analysis
Although MotiF shows clear advantages over prior work, it is still far from perfect for TI2V generation especially on our proposed challenging benchmark TI2V-Bench. As shown in Figure A3, we observe two main types of failure cases. First, sometimes the generated motion is not very natural. Second, the generated video may not follow the text prompt. In the second case, there are two challenging scenarios of TI2V-Bench: 1) when the text prompt describes a new object that needs to appear in the scene, the generated video may not be very coherent; 2) when there are multiple objects and the text prompt only refers to one of them, it will be hard for the model to generate precise motion.
We hope MotiF and TI2V-Bench will help the research community to tackle this challenging problem. MotiF is generic and complementary to existing techniques for TI2V generation. We believe that improving the motion heatmap accuracy can potentially boost the performance. Moreover, MotiF is potentially applicable to text-to-video generation.
@@ Appendix D Additional Implementation Details
D.1 Computational Costs of Optical Flow
The average speed of optical flow generation is 2.1 videos per second on a single Nvidia A100 GPU. The speed remains the same for videos of higher resolution, as the model resizes all video frames to a fixed resolution of 960 × 520.
D.2 Scenes in TI2V-Bench
TI2V-Bench contains a total of 22 diverse scenes, designed to cover a wide range of scenarios. The detailed scenes include: car on the road, balance scale, (multiple) balloons, bird, (multiple) bulbs, butterfly, candle, child in a playground, dog, rubber duck on a pool, fish, flower, golf ball, horse, animal on a meadow, human face, human body, sun, tide, traffic light, tree, and volcano.
D.3 Human Evaluation
We show the human evaluation interface in A4. During the evaluation process, the annotators are required to read the instructions first and then answer the two questions based on the specified criteria.